MONT 105Q – Mathematical Journeys Second Writing Assignment April, 2016

The second paper assignment this semester will be devoted to topics connected with our final CHQ common text, the play Hamlet by William Shakespeare.

For this assignment you have a choice between a more serious essay and a somewhat lighter topic with a chance to show more creativity.

Topic 1 – The more serious essay

In Act V, scene 1, 285-287 Hamlet says to Laertes beside the grave of Ophelia: “forty thousand brothers could not with all their quantity of love make up my sum.” In all the writing that has been done about Hamlet, the ways the characters feel and express the emotion of love is much less often discussed than characters’ desires for revenge or Hamlet’s (real or feigned) madness. Yet love plays a key role in the tragedy. Considering Hamlet, Gertrude, Claudius, Polonius, Laertes, and Ophelia in turn:

• identify who (or what) it is that they really love (you may want to claim that this is different from what they say at times; provide evidence chosen from the text of the play to make your points)

• determine whether that love is returned or not (a liberal dose of interpretation will be necessary in some cases!),

• decide how or whether each character is changed by who or what they have loved and how that love was returned or not.

Assemble all your analysis into a “shapely” essay – that is, you should not end up with just six paragraphs, one on each of the characters(!)

Topic 2 – The lighter one

Imagine that you are a reporter for the celebrity gossip webpage TMZ.com. You have been assigned to cover the recent bloodbath at the castle of Elsinore described in the play Hamlet (but assuming the action has taken place in the near past of 2016). Your editor has given you a 3-5 page limit for the “copy” (the text) for your post (someone else will create the web page with racy pictures of blood-covered bodies). This is longer than the usual story on the page, but that is justified since so much happened(!) Your readers want to find out exactly what happened (with all the “juicy details” like the story of a ghostly apparition, the visit of the traveling players and their production of the Murder of Gonzago, the drowning of Ophelia, etc.). They also want enough of the “backstory” to feel that they understand why it all happened as it did. A blow-by-blow account of the whole sequence of events (that is, the whole plot of the play) would be too long and too boring to satisfy anyone. So you will have to think about how to organize the story, and you will need to make some choices about what to include. Summarize appropriately and keep it lively and “catchy” for the typical ADHD TMZ.com readership(!)

Whichever topic you choose, your writing should show that you have done some careful reading and some serious thinking about the play. You may, of course, disagree with ideas presented in class if you want. The point as always is for you to develop your own thinking, and then clearly express it in writing.

Your paper should be 3-5 pages in length (double-spaced, single-sided). Please submit as a MS Word .doc or .docx file to jlittle@holycross.edu by 5:00pm on Friday, April 29. (Plan carefully so that work on this assignment does not interfere with your work on the final project!)

Guidelines for this assignment

Follow these specifications:

1. It is OK to discuss preliminary ideas for what you want to say with your classmates if you want. I will also be happy to talk over things and/or read a preliminary draft to give feedback if you like.

2. You may consult other sources beyond Hamlet itself if you want.

3. If you use direct quotations, identify the source in a foot- or end-note and provide a full citation.

4. The paper as a whole should be well-organized and “flow” from one point to the next. Your sentences should be clear, readable, and not overly wordy.

5. Especially for the second topic, this paper can be less formal. You should aim for a lively style and the writing should be in your own voice. Very colloquial ways of saying things are OK if they contribute to the point you are trying to make. But do it for a reason.

6. Please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

7. Errors in mechanics and usage should not annoy the reader. (As a warning, I tend to be easily annoyed by this sort of thing!) Try as much as possible to avoid mistakes in grammar, spelling, punctuation, and word choice.